OpenStudy (anonymous):

BEST ANSWER What is the 32nd term of the arithmetic sequence where a1 = -34 and a9 = -122? -408 -397 -386 -375

OpenStudy (anonymous):

Since we are looking at an arithmetic sequence, each term is the sum of the previous term and sum number (which we will call 'd'). Since a1 = -34 and a9 = -122, you can solve for 'd' (also known as the 'difference') by doing: \[-122 - (-34) = -88\] \[-88/8 = -11 = d\] As I mentioned before, 'd' is the 'difference' between each term in the arithmetic sequence. Using the general formula for an arithmetic sequence, aN = a1 + d(N-1) (where 'N' is the number of the term you want in the sequence), you can find the 32nd term.